OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. C(0, 4), m = 0

OpenStudy (anonymous):

@justpicksomethin help!!

OpenStudy (anonymous):

y-4=0(x-0)

OpenStudy (anonymous):

^ that how u do it :o

OpenStudy (anonymous):

hahah did you even read the lesson?

OpenStudy (anonymous):

i did.. im 4 weeks behind

OpenStudy (anonymous):

ive read it 4 weeks -.-

OpenStudy (alexwee123):

y-4=0

OpenStudy (anonymous):

 @jollysailorbold set it up corectly